# ManiGaussian: Dynamic Gaussian Splatting for Multi-task Robotic Manipulation

### v1

@@ 3.4 Learning Objectives
where and respectively mean the prediction and groundtruth of observation images from different views at the step.
@@
where , , , represent the probability of the groundtruth actions for translation, rotation, gripper openness and motion planner states of the robot, respectively.
@@
where means the predicted future observation of the scene at the step based on the action and current observation , and is the realistic counterpart. The predicted action at the step is acquired from the action decoder according to the Gaussian parameters, and the reconstructed observation for current scene can be obtained by rendering the current Gaussian parameters. The objective enforces the dynamic Gaussian Splatting framework to learn informative Gaussian parameters to generate optimal actions and recover the current observation, so that the future observation can be accurately reconstructed with embedded spatiotemporal dynamics.
@@
where are the hyperparamters that controls the importance of different terms during training. In training, we set a warm-up phase that freezes the deformation predictor to first learn a stable representation model and a Gaussian regressor during the first k iterations. After the warm-up phase, we jointly train the whole Gaussian world model with the action decoder until the training ends.
@@ 4.1 Experiment Setup
Dataset. Our experiments are conducted in the popular RLBench [24] simulated tasks. Following [69], we utilize a curated subset of challenging language-conditioned manipulation tasks from RLBench, which includes variations in object properties and scene arrangement. The diversity of these tasks requires the agent to acquire generalizable knowledge about the intrinsical scene-level spatial-temporal dynamics for manipulation, rather than solely relying on mimicking the provided expert demonstrations to achieve high success rates. We evaluated 25 episodes in the testing set for each task to avoid result bias from noise. For visual observation, we employ RGB-D images captured by a single front camera with a resolution of . During the training phase, we use demonstrations collected by GNFactor [69] for each task.
@@
Implementation Details. We use the [55, 69] augmentation for the expert demonstrations in the training set to enhance the generalizability of agents. To ensure consistency and mitigate the impact of parameter size, we utilize the same version of PerceiverIO [23] as the action decoder across all baselines. All the compared methods are trained on two NVIDIA RTX 4090 GPUs for k iterations with a batch size of . We employ LAMB optimizer [67] with the initial learning rate . We also adopt a cosine scheduler with warmup in the first 3k steps.
| Method / Task |
close jar |
open drawer |
sweep to dustpan |
meat off grill |
turn tap |
@@
4.2 Comparison with the State-of-the-Art Methods.
In this section, we compare our ManiGaussian with previous state-of-the-art methods on the RLBench tasksuite. Table 1 illustrates the comparison of the average success rate of each task. Our method achieves the best performance with an average success rate of %, which is the state-of-the-art, outperforming the previous arts including both perceptive and generative-based methods by a sizable margin. The dominated generative-based method GNFactor [69] leveraged a generalizable NeRF to learn informative latent representation for optimal action prediction, which showed effective improvement beyond the perceptive-based methods. However, it ignores the scene-level spatiotemporal dynamics that demonstrate the interaction among objects, and the predicted actions still fail to achieve human goals because of the incorrect interaction. On the contrary, our ManiGaussian learns the scene dynamics with the proposed dynamic Gaussian Splatting framework, so that the robotic agent can complete human instructions with accurate action prediction in unstructured environments. As a result, our method outperforms the second-best GNFactor method by a relative improvement of %. In addition, out of tasks, we excel in tasks, achieving more than double the accuracy of the previous methods in the “sweep to dustpan”, “put in drawer” and “drag stick” tasks. In the task “meat off grill” where the best performance was not reached, our method also ranks as second best. In conclusion, the experimental results illustrate the effectiveness of our proposed method across multiple language-conditioned robotic manipulation tasks.
@@ 4.3 Ablation Study
Our dynamic Gaussian Splatting framework models the propagation of diverse features in the Gaussian embedding space, and the Gaussian world model reconstructs the future scene according to the current scene by constraining the consistency between reconstructed and realistic scenes for dynamics mining. We conduct an ablation study to verify the effectiveness of each presented component in Table 2, where we manually group the RLBench tasks into categories according to their main challenges to demonstrate the improvement reason. We first implement a vanilla baseline without any proposed technique, where we directly train the representation model and the action decoder in the Gaussian world model to predict the optimal robot action. By adding the Gaussian regressor to predict the Gaussian parameters in the dynamic Gaussian Splatting framework, the performance improves by % compared with the vanilla baseline. Especially, in the tasks that require geometric reasoning such as “Occulusion”, “Tools” and “Screw”, it outperforms the vanilla version by sizable margins, which proves the ability of the Gaussian Splatting technology to model the spatial information for manipulation tasks. We then add semantic features distilled from the pretrained foundation model into the dynamic Gaussian Splatting framework, which contains high-level visual information of the observed scenes. By adding the semantic features and the related consistency loss, we observe that the average success rate increases by % than the only geometric features version, which indicates the benefits of the rich semantic information for robotic manipulation. Besides, we implement the deformation predictor and the corresponding future scene consistency loss based on the vanilla baseline to mine the scene-level dynamics via future scene reconstruction, resulting in a dramatic performance improvement of %. Particularly, the proposed deformation predictor improves the task completion of out of task types, including “Planning”, “Long”, “Motion”, and “Occlusion”, which demonstrates the importance of the scene-level dynamics encoded by the deformation predictor in the Gaussian world model. After combining all the proposed techniques in our dynamic Gaussian Splatting framework, the performance increases from % to %, which verifies the necessity of the scene-level spatiotemporal dynamics mined by the proposed dynamics Gaussian Splatting framework with the Gaussian world model parameterization.
Figure 3 shows the learning curve of the proposed ManiGaussian and the state-of-the-art method GNFactor, where we save checkpoints and test them every k parameter updates. Both the compared methods get convergence within k training steps. As shown in Figure 3, our ManiGaussian outperforms the state-of-the-art method GNFactor, achieving better performance and faster training. This result proves that our ManiGaussian not only performs better but also trains faster, which also shows the efficiency of the explicit Gaussian scene reconstruction than the implicit approach like NeRF.
@@ 4.4 Qualitative Analysis
Visualization of Whole Trajectories. We present two qualitative examples of the generated action sequence in Figure 4 from GNFactor and our ManiGaussian. In the top case, the agent is instructed to “Slide the block to yellow target”. The results show that the previous agent struggles to complete the task since It imitates the expert’s backward pulling motion, even though the claw is already leaning towards the right side of the red block. In contrast, our ManiGaussian returns to the red square and successfully slides the square to the yellow target, owing to that our method can correctly understand the scene dynamics of objects in contact. In the bottom case, the agent is instructed to “Turn left tap”. The results show that the GNFactor agent misunderstands the meaning of “left”, and instead operates the right tap, and also fails to turn on the tap. Our ManiGaussian successfully completes the task, which shows that our method can not only understand the semantic information, but also execute operations accurately.
Visualization of Novel View Synthesis. Figure 5 shows the novel view image synthesis results, where the action loss is removed for all compared methods for better visualization. The results show that, first, based on the front view observation where the claw shape cannot be seen, our ManiGaussian offers superior detail in modeling cubes in novel views within the same number of training steps. Second, our method accurately predicts future states and can model interactions between objects based on accurately recovering object details. For example, in the top case of the “slide block” task, our ManiGaussian not only predicts the future gripper position that corresponds to the human instruction, but also predicts the future object location influenced by the gripper based on the scene dynamics understanding of the physical interaction among objects. This qualitative result demonstrates that our ManiGaussian learns the intricate scene-level dynamics successfully.
@@ 5 Conclusion
In this paper, we have presented a ManiGaussian agent that encodes the scene-level spatiotemporal dynamics for language-conditioned manipulation agents. We design a dynamic Gaussian Splatting framework that models the propagation of diverse semantic features in the Gaussian embedding space, and the semantic features with scene dynamics are leveraged to predict the optimal robot actions. Subsequently, we build a Gaussian world model to parameterize the distributions in the dynamic Gaussian Splatting framework to mine scene-level dynamics by reconstructing the future scene according to the current scene and the robot actions. Extensive experiments in a wide variety of manipulation tasks demonstrate the superiority of our ManiGaussian compared with the state-of-the-art methods. The limitations of our ManiGaussian stem from the necessity of multiple view supervision with camera calibration for the dynamic Gaussian Splatting framework.

### v2

@@ 3.4 Learning Objectives
where and respectively mean the groundtruth and prediction of observation images from different views at the step.
@@
where , , , represent the probability of the groundtruth actions in expert demonstrations for translation, rotation, gripper openness and collision avoidance of the robot, respectively.
@@
where means the predicted future image of the scene at the step based on the action and current observation , and is the realistic counterpart. By imposing the Gaussian world model to predict future scenes based on the representation, the representation is required to encode the physical properties of the scene. This is important for the action decoder to predict effective actions with such representation.
@@
where are the hyperparamters that controls the importance of different terms during training. In training, we set a warm-up phase that freezes the deformation predictor to learn a stable representation model and a Gaussian regressor during the first k iterations. After the warm-up phase, we jointly train the whole Gaussian world model with the action decoder.
@@ 4.1 Experiment Setup
Simulation. Our experiments are conducted in the popular RLBench [27] simulated tasks. Following [76], we utilize a curated subset of challenging language-conditioned manipulation tasks from RLBench, which includes variations in object properties and scene arrangement. The diversity of these tasks requires the agent to acquire generalizable knowledge about the intrinsical scene-level spatial-temporal dynamics for manipulation, rather than solely relying on mimicking the provided expert demonstrations to achieve high success rates. We evaluated 25 episodes in the testing set for each task to avoid result bias from noise. For visual observation, we employ RGB-D images captured by a single front camera with a resolution of . We use the same number of cameras (i.e., ) as GNFactor to provide multi-view supervision for fair comparisons. During the training phase, we use demonstrations for each task.
@@
Implementation Details. We use the [60, 76] augmentation for the expert demonstrations in the training set to enhance the generalizability of agents. To mitigate the impact of parameter size, we utilize the same version of PerceiverIO [26] as the action decoder across all baselines. All the compared methods are trained on two NVIDIA RTX 4090 GPUs for k iterations with a batch size of . We employ LAMB optimizer [74] with an initial learning rate . We also adopt a cosine scheduler with warmup in the first 3k steps.
| Method / Task | close jar | open drawer | sweep to dustpan | meat off grill | turn tap |
@@
4.2 Comparison with the State-of-the-Art Methods
In this section, we compare our ManiGaussian with previous state-of-the-art methods on the RLBench tasksuite. Table 1 illustrates the comparison of the average success rate of each task. Our method achieves the best performance with an average success rate of %, which is state-of-the-art, outperforming the previous arts including both perceptive and generative-based methods by a sizable margin. The dominated generative-based method GNFactor leveraged a generalizable NeRF to learn informative latent representation for optimal action prediction, which showed effective improvement beyond the perceptive-based method PerAct. However, it ignores the scene-level spatiotemporal dynamics that demonstrate the interaction among objects, and the predicted actions still fail to achieve human goals because of the incorrect interaction. On the contrary, our ManiGaussian learns the scene dynamics with the proposed dynamic Gaussian Splatting framework, so that the robotic agent can complete human instructions with accurate action prediction in unstructured environments. As a result, our method outperforms the second-best GNFactor method by a relative improvement of %. In the task meat off grill where the best performance was not reached, our method also ranks as second best. The experimental results illustrate the effectiveness of our proposed method across multiple language-conditioned robotic manipulation tasks.
@@ 4.3 Ablation Study
Our dynamic Gaussian Splatting framework models the propagation of diverse features in the Gaussian embedding space, and the Gaussian world model reconstructs the future scene according to the current scene by constraining the consistency between reconstructed and realistic scenes for dynamics mining. We conduct an ablation study to verify the effectiveness of each presented component in Table 2. We first implement a vanilla baseline without any proposed technique, where we directly train the representation model and the action decoder to predict the robot actions. By adding the Gaussian regressor to predict the Gaussian parameters, the performance improves by % compared with the baseline.
Especially, in the tasks that require geometric reasoning such as Occlusion, Tools and Screw, it outperforms the vanilla version by sizable margins, which proves the ability of the Gaussian Splatting technology to model the spatial information for manipulation tasks. We then add semantic features distilled from the pretrained foundation model into the dynamic Gaussian Splatting framework. By adding the semantic features and the related consistency loss, we observe that the average success rate increases by % than the only geometric features version, which indicates the benefits of the high-level semantic information for robotic manipulation. Besides, we implement the deformation predictor and the corresponding future scene consistency loss, resulting in a dramatic performance improvement of %. Particularly, the proposed deformation predictor improves the task completion of out of task types, which demonstrates the importance of the scene-level dynamics encoded by the deformation predictor in the Gaussian world model, especially in long-horizon tasks (Long). Though the dynamic loss may slightly impact short-term results due to the balance of different loss items, it significantly improves overall performance. After combining all the techniques in our dynamic Gaussian Splatting framework, the performance increases from % to %, which verifies the necessity of the scene-level spatiotemporal dynamics mined by the proposed dynamics Gaussian Splatting framework with the Gaussian world model.
Figure 3 depicts the learning curve of the proposed ManiGaussian and the state-of-the-art method GNFactor, where we save checkpoints and test them every k parameter updates. Both the compared methods get convergence within k training steps. As shown in Figure 3, our ManiGaussian outperforms the state-of-the-art method GNFactor, achieving better performance and faster training. This result proves that our ManiGaussian not only performs better but also trains faster, which also shows the efficiency of the explicit Gaussian scene reconstruction than the implicit approach like NeRF.
@@ 4.4 Qualitative Analysis
Visualization of Whole Trajectories. We present two qualitative examples of the generated action sequence in Figure 4 from GNFactor and our ManiGaussian. In the top case, the agent is instructed to “Slide the block to yellow target”. The results show that the previous agent struggles to complete the task since it imitates the expert’s backward pulling motion, even though the claw is already leaning towards the right side of the red block. In contrast, ManiGaussian returns to the red square and successfully slides the square to the yellow target, owing to that our method can correctly understand the scene dynamics of objects in contact. In the bottom case, the agent is instructed to “Turn left tap”. The results show that GNFactor misunderstands the meaning of “left”, and instead operates the right tap, and also fails to turn on the tap. In contrast, our ManiGaussian successfully completes the task, which shows that ManiGaussian can not only understand the semantic information, but also execute operations accurately.
Visualization of Novel View Synthesis. Figure 5 shows the novel view image synthesis results. First, based on the front view observation where the gripper shape cannot be seen, our ManiGaussian offers superior detail in modeling cubes in novel views. Second, our method accurately predicts future states based on the recovered details. For example, in the top case of the slide block task, our ManiGaussian not only predicts the future gripper position that corresponds to the human instruction, but also predicts the future cube location influenced by the gripper based on the understanding of the physical interaction among objects. This qualitative result demonstrates that our ManiGaussian learns the intricate scene-level dynamics successfully.
@@ 5 Conclusion
In this paper, we have presented a ManiGaussian agent that encodes the scene-level spatiotemporal dynamics for language-conditioned manipulation agents. We design a dynamic Gaussian Splatting framework that models the propagation of features in the Gaussian embedding space, and the latent representation with scene dynamics is leveraged to predict the robot actions. Subsequently, we build a Gaussian world model to parameterize the distributions in the dynamic Gaussian Splatting framework to mine scene-level dynamics by reconstructing the future scene. Experiments in diverse manipulation tasks demonstrate the superiority of ManiGaussian. The limitations stem from the necessity of multiple view supervision with camera calibration for the Gaussian Splatting framework.
Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under Grant 2022ZD0114903 and Shenzhen Ubiquitous Data Enabling Key Lab under grant ZDSYS20220527171406015. We would like to greatly thank Yanjie Ze for his kind response to the GNFactor issues.
Supplementary Material
In this supplementary material, we provide additional details and experiments not included in the main paper due to limitations in space.
Appendix 0.A: Details of the RLBench dataset and the training pipeline used in our experiments.
Appendix 0.B: Additional implementation details of our ManiGaussian.
Appendix 0.C: Supplementary quantitative analysis.
Appendix 0.D: Supplementary qualitative analysis.
Appendix 0.A Details of RLBench
RLBench Dataset. In this section, we provide a concise overview of the RLBench [27] dataset and our training pipeline. Table 3 is an overview of the selected tasks we use in the experiments. Our task variations include randomly sampled colors, sizes, counts, placements, and categories of objects. We have a color palette of shades, including red, maroon, lime, green, blue, navy, yellow, cyan, magenta, silver, gray, orange, olive, purple, teal, azure, violet, rose, black, and white. The size of the objects is categorized into two types: short and tall. The number of objects can be either , , or . Other properties vary depending on the specific task. Furthermore, objects are randomly arranged on the tabletop within a certain range, adding to the diversity of the tasks. In the ablation study, we adopt the task classification from [15] to group the RLBench tasks of Table 3 into categories according to their key challenges. The task groups include:
The Planning group contains tasks with multiple subtasks. The included tasks are: meat off grill and push buttons.
The Long group includes long-term tasks that requires more than keyframes. The included tasks are: put in drawer and stack blocks.
The Tools group requires the agent to grasp an object to interact with the target object. The included tasks are: slide block, drag stick and sweep to dustpan.
The Motion group requires precise control, which often causes failures due to the predefined motion planner. The included task is: turn tap.
The Screw group requires gripper rotation to screw an object. The included task is: close jar.
The Occlusion group involves tasks with severe occlusion problems from certain views. The included task is: open drawer.
| Task | Type | Variations | Keyframes | Instruction Template |
| close jar | color | 20 | 6.0 | “close the jar” |
| open drawer | placement | 3 | 3.0 | “open the drawer” |
| sweep to dustpan | size | 2 | 4.6 | “sweep dirt to the dustpan” |
| meat off grill | category | 2 | 5.0 | “take the off the grill” |
| turn tap | placement | 2 | 2.0 | “turn tap” |
| slide block | color | 4 | 4.7 | “slide the block to target” |
| put in drawer | placement | 3 | 12.0 | “put the item in the drawer” |
| drag stick | color | 20 | 6.0 | “use the stick to drag the cube onto the target” |
| push buttons | color | 50 | 3.8 | “push the button, [then the button]” |
| stack blocks | color, count | 60 | 14.6 | “stack blocks” |
Training Pipeline. To learn the policy, we uniformly sample a group of expert episodes from all the task variations, and then randomly choose an input-action pair for each of the tasks to form a batch. Other sampling strategies (e.g., Auto- [41]) are also available. To simplify the tasks, the agent is assumed to access a predefined motion planner (e.g., RRT-Connect), so that the input-action pairs are determined as the bottleneck end-effector poses (i.e., keyframes) within each demonstration based on empirical rules: a pose is determined as a keyframe if the end-effector changes state (e.g. close the gripper) or its velocities approach zero [7, 13, 28, 60, 76]. This setting simplifies the sequential decision-making problem into predicting the next optimal keyframe action based on the current observation, which can also be interpreted as a classification task.
Appendix 0.B Additional Implementation Details
In this section, we detail the architectural design of each submodule in our Gaussian world model. For more details, please refer to our code.
Representation model. The representation model is the same with [76], which is not the main contribution in this paper. The representation model utilize a shallow 3D UNet to encode the voxel (RGB features, coordinates, indices, and occupancy) into the high-level visual features .
Gaussian regressor. Given the current features encoded by the representation model , we pass it through a generalizable Gaussian regressor to infer the Gaussian distribution directly. The Gaussian regressor is designed as a lightweight multi-head neural network, where each head is responsible for predicting a specific feature. It consists of: (1) a position offset head that predicts the per-pixel 3D center offset , (2) a color head that predicts the coefficients of the spherical harmonic basis , (3) a rotation head with normalization that predicts the rotation quaternion , (4) a scaling head with exponential activation that outputs the scaling factor , (5) an opacity head with sigmoid activation that predicts the opacity . (6) a semantic head that predicts the semantic feature .
Deformation predictor. After obtaining the current visual features , Gaussian embedding and action , we parameterize the transition process as a deformation predictor to predict the deformation and of each Gaussian, resulting in the future Gaussian embedding . The deformation predictor is a fully-connected network with residual connections.
Hyperparameters. The hyperparameters used in ManiGaussian are shown in Table 4. To train the robotic manipulation agent, we use , and to focus on the action prediction. Other hyperparameters are in line with previous works [60, 76] for fair comparison.
| Hyperparameter | Value |
| training iteration | k |
| image resolution | |
| voxel resolution | |
| batch size | |
| optimizer | LAMB |
| learning rate | |
| weight decay | |
| Number of Gaussian points | |
Appendix 0.C Additional Quantitative Analysis
We provide further ablation study on different implementation choices in our ManiGaussian. Table 5 presents the impact of different balance hyperparameters on the overall performance, from which we can conclude that the balance of each loss item is important to learn an optimal manipulation policy.
| Planning | Long | Tools | Motion | Screw | Occlusion | Average | |||
| 42.0 | 24.0 | 48.0 | 48.0 | 28.0 | 72.0 | ||||
| 54.0 | 12.0 | 44.0 | 52.0 | 28.0 | 80.0 | ||||
| 43.6 | |||||||||
| 48.0 | 8.0 | 34.7 | 48.0 | 24.0 | 64.0 | ||||
| 41.6 | |||||||||
| 46.0 | 2.0 | 37.3 | 60.0 | 40.0 | 68.0 | ||||
| 44.8 |
Appendix 0.D Additional Qualitative Analysis
We provide additional comprehensive episodes generated by our ManiGaussian and the state-of-the-art generative method GNFactor [76] in the attached video file (demo.mp4). In the long-term “stack 2 rose blocks”, “put the item in the bottom drawer” and “take the steak off the grill” tasks, ManiGaussian completes the human instructions in the correct order with the understanding of high-level scene dynamics mined by the Gaussian world model. In the “sweep dirt to the short dustpan” and “use the stick to drag the cube onto the azure target” tasks that involve tool usage, our ManiGaussian succeeds in solving the tasks by correctly understanding the low-level scene dynamics of objects in contact. In the “slide the block to green target”, “turn left tap”, “close the azure jar” and “open the bottom drawer” tasks that require semantic understanding and precise control, our ManiGaussian can successfully comprehend the semantic information to interact with the correct object instance, while the baseline method often confuses different instances.